Find the value of k, if x = 2, y = 1 is a solution of the equation 2x + 3y = k.

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear student

The given equation is

2x + 3y = k

According to the question, x = 2 and y = 1.

Substituting the values of x and y in the equation 2x + 3y = k,

We get,

⇒(2 x 2)+ (3 × 1) = k

⇒4+3 = k

⇒7 = k

⇒k = 7

The value of k is 7.
Thanks